# Two dimensional de-Sitter and deformed CFTs

### v1

@@ 1 Introduction
Lately, there has been considerable work [12]-[24] [^2] This is not exhaustive, see the cited works for more. exploring the idea of including the observer as part of the gravitational system in a closed universe following [11]. The most natural setting to investigate this proposal is a closed de Sitter space-time, where a static observer’s accessible region is confined to a static patch bounded by cosmological (as well as black hole) horizons. Apart from a theoretical playground, dS universe has a more direct contact with observational cosmology and the fate of our living universe. The widely studied inflationary regime of cosmology has been approximately described by a metastable dS phase. On the other hand, the present observation of accelerated expansion also suggests that the universe will go through an asymptotic de-sitter phase at late time. If we imagine an astronomical observer of the current time, it will probably see a locally dS static patch in his own clock after sufficiently long time. Hence understanding the huge amount of ‘mysterious’ Hawking-Gibbons entropy of static patch [25] and consequently providing a workable model of ‘static patch holography’ would be a great deal of interest. The most obvious notion may come from the stretched horizon holography, where in the absence of an asymptotic boundary, the vast amount of entropy could be located near the horizon[26],[27]. More intuitively, if one may imagine an observer moving along the timelike stretched horizon, he could effectively store all the degrees of freedom and measure the same horizon entropy [^3] Note that, an interesting feature of the static patch cosmological horizon is it’s observer dependence. Even though, different observer could see different horizon, the area is always be the same.. Nevertheless, only a clean UV theory of static patch effective theory could lead us to the correct and proper understanding of static patch entropy and holography.
@@
By definition, the UV theory should have a low energy limit that describes (semi)classical theory of gravity [^4] This should be a minimal requirement for a consistent UV theory.. In the closed universe, as we mentioned, the UV theory should not have a definite notion of time. Here, in a modest sense, our goal of this note is to find a sensible theory that only describes an observer in the lower dimensional dS static patch in low energy limit. To be more precise, we are interested to study two dimensional effective theory of dS in specific coordinates and it’s possible UV description. With this motivation, we study S-wave dimensional reduction of asymptotic Schwarzchild dS4 geometry in the near nariai near horizon limit to find an effective 2D description of dilatons in section (2). In contrast to JT gravity, we provide an alternative reduction, in which the effective theory can be recast as free massless dilatons propagating in a static patch background enclosed by cosmological and blackhole horizon with same temperature. We quantize the dilatons using some UV dirichlet boundary condition at the stretched horizon and find a finite entropy which we interpret as out-side entropy $S_{out}$ of observers living inside the static patch with a worldline along the stretched horizon. In section (3), we consider S-wave reduction of pure dS3 in 3D Einstein gravity. Unlike a free theory, this effective theory(from any higher dimensional reduction) can be written as dynamical dilatons inside a past Milne wedge. We also observe the dilaton effective theory has an initial potential dominance and late time kinetic dominance, providing us a simple toy model for inflation in 2D. We relate it to a critical quench problem and then discuss a boundary non-preserving deformed CFT Hamiltonian on a strip in section 2.1. The ultimate fate of the constructed past Milne vacuum is shown to be unstable at (finite but) late time, where the kinetic term dominates over potential term in section (4.2). Since Milne patch is analytically continued to static patch with thermal Euclidean time circle, we immediately constructs a sensible worldsheet description of (extended) static patch [^5] Throughout the note, we refer ‘extended static patch’ as the dimensionally reduced Nariai black hole enclosed by black hole and cosmological horizon at same temperature. in terms of deformed CFT Hamiltonian on a ring in section (4). We further study it’s modular quantization and connect the thermal entropy with entanglement entropy of static patch vacuum in section (5.1) and (5.2). The boundary condition on the stretched horizon is an emergent requirement of this quantization, as we also noticed in AdS [28],[1]. We also discuss our proposed alternative S-wave reduction in 4D magnetically charged RN AdS4 blackhole in the near horizon near extremal limit and find free massless dilatons coupled to boundary Schwarzian as an effective description in appendix (A). This provides a closure to our claim in a previous paper [1] on the UV completion of JT coupled to $c=\mathcal{O}(1)$ CFT as a worldsheet deformed CFT in a more concrete way.
@@
However, in this modest approach, the questions regarding the observer dependent choice of UV time is still obscure, since our description of worldsheet CFT already knows the ‘time’ of what observer will see in a static patch. We know in dS, different patches describe different notion of time and hence finding a global vacuum is inherently a difficult problem [^6] See [10] and the discussion in [21] from algebra of observables point of view.. We end our note with some comments on it in section 5 along with possible future direction.
@@ 2 Dimensional Reduction to 2D de-Sitter in near-Nariai limit
Here we review the dimensional reduction of pure Einstein gravity in 4D to two-dimensional JT gravity in the near-Nariai limit following [29]. Consider the following metric ansatz of the class of spherically symmetric asymptotically 4D de Sitter metric:
@@ 2.1 Alternative dual reduction in the near-Nariai limit
Hence, the entropy is $S=S_{N}+\log(A)$. This suggests, the total entropy may have some correction which is not observed from previous way. However, this is not a useful way to study dynamics of the effective theory. This motivates us to consider an inequivalent or alternative dimensional reduction which may capture another useful and tractable description of near horizon physics in closed universe.
@@ 2.2 Partition function of massless free scalar fields in a static patch background with stretched horizon
In section (5) we will explicitly compute and will try to match this entropy using the quantization of our proposed UV free bosonic deformed CFT on a ring. We note that, the Hamiltonian of dilatons with stretched horizon boundary condition defined by some UV theory, should be identified with the total Hamiltonian of the observer plus dilatons and the observer clock is set by the ‘time’ generated by static patch Hamiltonian due to Hamiltonian constraint.
@@ 5.1 Deformed CFT on a ring, emergent Virasoro algebra and stretched horizon in static patch
Using $z=\coth{\gamma\omega}$ and $\bar{z}=-\coth{\gamma\bar{\omega}}$, with $\omega=\theta+it^{\prime}$$\bar{\omega}=\theta-it^{\prime}$ and the transformation of stress tensor $T(\omega)=(\frac{\partial z}{\partial\omega})^{2}T(z)+\frac{c}{12}Sch\{z,\omega\}$, we end up with:
@@
Now $T(\omega)=\left[\frac{1}{\sinh^{4}{\gamma\omega}}T(z)+\frac{c\gamma^{2}}{12\pi}\right]$ and $\bar{T}(\bar{\omega})=\left[\frac{1}{\sinh^{4}{\gamma\bar{\omega}}}T(z)+\frac{c\gamma^{2}}{12\pi}\right]$. Since $\theta\rightarrow\pm\infty$ corresponds to $z\rightarrow\pm 1$, $T(z)|_{z=\pm 1}$, $\bar{T}(\bar{z})|_{\bar{z}=\pm 1}$ is finite. Thus we can immediately see [^25] Here again the stress tensor is negative but finite. Even though, one can have thermal two point function and other thermal features of the static patch, this constant negative stress tensor is not concerning. Rather the difference from an positive stress tensor in the AdS2 black hole vacuum is quite encouraging in the sense of distinguished nature of a cosmological horizon and a black hole horizon. An observer inside the cosmological horizon can effectively measure negative local energy density. However this static patch vacuum is not a (full) de-Sitter invariant vacuum. Time translation invariance in an expanding universe is just an artifact of forming causal horizons enclosing the observers.
@@ 5.2 Modular quantization on static patch and thermal partition function
  • We can construct eigenstates of regulated Hamiltonian $(\mathcal{L}_{0}^{\Lambda_{c}}+\bar{\mathcal{L}}_{0}^{\Lambda_{c}})$ which are localized at the cut-off surface or stretched horizon. These eigenstates are constructed out of primaries as well as modular Virasoro descendants. We note that, these states are emergent due to the emergent modular Virasoro algebra and the conformal boundary condition at the cut-off. All the states at the cut-off surface can be constructed explicitly for massless free bosons as states of regulated Fock space, briefly described in section 2.
  • Using this quantization, the $\epsilon\rightarrow 0$ limit can be identified with the zero modes of the Hamiltonian as we will show now. However, mapping back to $(\omega,\bar{\omega})$ plane these primaries are vanishing, leaving only vacuum at the horizon [^27] We took a different claim for these than our earlier paper [1], which we think is wrong.. Hence the vanishing edge modes should further give a vanishing micorcanonical entropy count sourced purely from the surviving vacuum state [^28] One can not obtain this just naively putting $\epsilon=0$ for which it is diverging. This different results further suggests that measuring entropy without a cut-off with boundary condition is meaningless.. This on the other hand tells, the entropy explicitly measures the stretched horizon observer entropy, which is not obtained in the semiclassical limit without an observer. The same applies in AdS, where one must impose a cutoff on the dilaton, which can be made arbitrarily large but cannot be taken literally to infinity.
@@
Hence the leading term matches with the entanglement entropy in a full time slice of the static patch $\theta:(-\Lambda_{c},\Lambda_{c})$ as one can check by computing twist operator two point function using the map $z\rightarrow\omega$. This is again consistent with the fact that the time translation of static patch Hamiltonian is identified with the flow by modular Hamiltonian on the full patch bounded by the horizons.
On the other hand, we can compare this result by explicit computation of free massless bosons($c=1$) in the stretched horizon background as in (2.40). We have seen a potential mismatch of overall factor of $\frac{1}{8}$. If we have 8 massless scalars in the static patch background, this would exactly match with a single scalar in the worldsheet description. Currently we have no clear understanding behind this mismatch. Neglecting that potential mismatch, we can see both of these terms proportional to $\log(\frac{2}{\epsilon})$ [^29] Here we took $r_{N}=1/\nu=1$. However for the general $r_{N}\neq 0$, we may need to consider $H_{c}=\alpha L_{0}+i\gamma(L_{1}-L_{-1})$ with $d=\alpha^{2}-4\gamma^{2}<0$. This class of Hamiltonian belongs to the ‘heating phase’ in the Floquet CFT language [65],[60]..
We end this section by commenting our take on the questions we have tried to address iff we may think of this deformed CFT quantization as an UV effect:
@@
    Thermal entropy of free massless bosons in stretched horizon background. Along the line of [60], one can again think the same as the entanglement entropy of the causal patch vacuum with the outside.  [^30] This is exactly the other side of the quantization contour as described in [60]..
@@
    What is the static patch holography in this context?
    All the entropic degrees of freedom of the extended static patch is restored in the stretched horizon and the Hilbert space is defined by the modular quantization.
@@
    The stretched horizon or the moving membrane defined by the worldsheet deformed CFT on a ring.
@@ 6 Discussion
In this note, our main goal is to find possible UV description of effective 2D theory of dilatons arising from reduction of higher dimensional Einstein gravity describing pure and asymptotically dS. While a stable dS2 vacuum is difficult to find in any time-dependent patches of dS2, we have managed to describe a sensible description of (extended and not pure) static patch effective theory in terms of worldsheet CFT with deformed Hamiltonian. Modular quantization of the worldsheet CFT further introduces a stretched horizon with an emergent conformal boundary condition, which acts like an internal observer in this description. It would be definitely interesting to find a set-up where negatively divergent energy density in Milne vacuum could be compensated by gluing it with a very high positive energy density state like large AdS2 blackhole. This could probably lead to some version of cosmology in open universe.
@@
[…]
We end our discussion with some speculative remarks on the emergence of worldsheet time. Serious readers may skip this part.
@@ 6.1 Comment on emergence of observer time
Nevertheless, the main point we want to make here is that the time $t$ inherent to the Floquet Hamiltonian $H_{F}$, is a stroboscopic time $t=n(T_{1}+T_{2})$. In this Floquet sense, the stroboscopic time is emergent. To be more precise, the Floquet observer will measure stroboscopic time in their own clock which is completely different than the clocks for the observers associated to $H_{1,2}$. This is analogous to find a static patch in an accelerating or time dependent (dS) universe. Consequently, to account for the cost of residing in an accelerating universe, static observers are required to observe features like surrounding cosmological (and black hole) horizons. We know that Floquet Hamiltonians in heating phases also exhibit features such as fixed points. We feel, the gravitational observers measure the stroboscopic Floquet time in the worldsheet description, which is inherently emergent. The principle behind the fine tuning of choosing the Hamiltonians $H_{1,2}$, drive parameters and frequencies still remains elusive to us from a gravitational perspective.
Acknowledgement: The author(SD) of this paper strongly believes that any scientific progress is a collective and collaborative process where the authorship is an effort to consolidate all the ideas in a single manuscript. He would like to thank Bobby Ezhuthachan, Rahul Roy and Krishnendu Sengupta for discussion at the early stage of this project. He also wants to thank collectively Dibya Chakraborty, Anirban Dinda, Arnab Kundu, Parthasarathi Majumdar, Debangshu Mukherjee and Koushik Ray, as well as ChatGPT and Perplexity for related discussions and help. He is also grateful to the co-authors of [60]: Bobby Ezhuthachan, Somnath Porey and Baishali Roy for noticing and emphasizing the paper by Solodukhin [42], which provides an important clue to understand the reduction process of this note. He would also like to acknowledge the organisers of the school “Introductory School on Conformal Field Theories” at IACS(18-22 Aug,2025), which helps him to clarify some basic 2D CFT stuff. The research work of SD is supported by a DST INSPIRE Faculty Fellowship.
@@
In [1], we claim that the classical limit of JT gravity coupled to CFT with $\mathcal{O}(1)$ central charge is described by a certain deformed CFT in a prescribed limit. Partial justification of this claim was made by the fact that in the semiclassical limit, the backreaction of the conformal fields can be neglected and the dynamics of JT coupled to matter CFT is boiled down to the dynamics of CFT in the UV frame. We recovered the entropy of the coupled system away from extremality, purely from the entropy of deformed CFT in the prescribed classical limit. Here in this appendix, we provide a concrete set-up for this justification where JT coupled to free bosons with $c=1$ arises naturally from S-wave reduction in the near extremal near horizon limit of 4D RN black holes in AdS [91].
@@
In either description, the Hilbert space of modular quantization or the Hilbert space of free scalars in stretched horizon background of AdS2 black hole provides the UV Hilbert space. Perhaps, the 3d version of this description has been realized in [36] [^44] I thank Chethan Krishnan for emphasizing this possibility at some point while discussing [36]..

### v2

@@ 1 Introduction
Lately, there has been considerable work [12]-[24] [^2] This is not exhaustive, see the cited works for more. exploring the idea of including the observer as part of the gravitational system in a closed universe following [11]. The most natural setting to investigate this proposal is a closed de Sitter space-time, where a static observer’s accessible region is confined to a static patch bounded by cosmological (as well as black hole) horizons. Apart from a theoretical playground, dS universe has a more direct contact with observational cosmology and the fate of our living universe. The widely studied inflationary regime of cosmology has been approximately described by a metastable dS phase. On the other hand, the present observation of accelerated expansion also suggests that the universe will go through an asymptotic de-sitter phase at late time. If we imagine an astronomical observer of the current time, it will probably see a locally dS static patch in his own clock after sufficiently long time. Hence understanding the huge amount of ‘mysterious’ Hawking-Gibbons entropy of static patch [25] and consequently providing a workable model of ‘static patch holography’ would be a great deal of interest. The most obvious notion may come from the stretched horizon holography, where in the absence of an asymptotic boundary, the vast amount of entropy could be located near the horizon[26],[27],[28]. More intuitively, from the perspective of a podal observer(located at the center of static patch), all the degrees of freedom may be captured by the timelike stretched horizon, where the stretched horizon itself plays the role of a gravitating observer [^3] Note that, an interesting feature of the static patch cosmological horizon is it’s observer dependence. Even though, different observer could see different horizon, the area is always be the same.. Nevertheless, only a clean UV theory of static patch effective theory could lead us to the correct and proper understanding of static patch entropy and holography.
@@
By definition, the UV theory should have a low energy limit that describes (semi)classical theory of gravity [^4] This should be a minimal requirement for a consistent UV theory.. In the closed universe, as we mentioned, the UV theory should not have a definite notion of time. Here, in a modest sense, our goal of this note is to find a sensible theory that only describes an observer in the lower dimensional dS static patch in low energy limit. To be more precise, we are interested to study two dimensional effective theory of dS in specific coordinates and it’s possible UV description. With this motivation, we study S-wave dimensional reduction of asymptotic Schwarzchild dS4 geometry in the near nariai near horizon limit to find an effective 2D description of dilatons in section (2). In contrast to JT gravity, we provide an alternative reduction, in which the effective theory can be recast as free massless dilatons propagating in a static patch background enclosed by cosmological and blackhole horizon with same temperature. We quantize the dilatons using some UV dirichlet boundary condition at the stretched horizon and find a finite entropy which we interpret as out-side entropy $S_{out}$ of observers living inside the static patch with a worldline along the stretched horizon. In section (3), we consider S-wave reduction of pure dS3 in 3D Einstein gravity. Unlike a free theory, this effective theory(from any higher dimensional reduction) can be written as dynamical dilatons inside a past Milne wedge [^5] Here past Milne wedge refers to the past of $I^{+}$ or the upper wedge in the Penrose diagram as in fig (1.b) of [29].. We also observe the dilaton effective theory has an initial potential dominance and late time kinetic dominance, providing us a simple toy model for inflation in 2D. We relate it to a critical quench problem and then discuss a boundary non-preserving deformed CFT Hamiltonian on a strip in section 2.1. The ultimate fate of the constructed past Milne vacuum is shown to be unstable at (finite but) late time, where the kinetic term dominates over potential term in section (4.2). Since Milne patch is analytically continued to static patch with thermal Euclidean time circle, we immediately constructs a sensible worldsheet description of (extended) static patch [^6] Throughout the note, we refer ‘extended static patch’ as the dimensionally reduced Nariai black hole enclosed by black hole and cosmological horizon at same temperature. in terms of deformed CFT Hamiltonian on a ring in section (4). We further study it’s modular quantization and connect the thermal entropy with entanglement entropy of static patch vacuum in section (5.1) and (5.2). The boundary condition on the stretched horizon is an emergent requirement of this quantization, as we also noticed in AdS [30],[1]. We also discuss our proposed alternative S-wave reduction in 4D magnetically charged RN AdS4 blackhole in the near horizon near extremal limit and find free massless dilatons coupled to boundary Schwarzian as an effective description in appendix (A). This provides a closure to our claim in a previous paper [1] on the UV completion of JT coupled to $c=\mathcal{O}(1)$ CFT as a worldsheet deformed CFT in a more concrete way.
@@
However, in this modest approach, the questions regarding the observer dependent choice of UV time is still obscure, since our description of worldsheet CFT already knows the ‘time’ of what observer will see in a static patch. We know in dS, different patches describe different notion of time and hence finding a global vacuum is inherently a difficult problem [^7] See [10] and the discussion in [21] from algebra of observables point of view. However, our present understanding might be more aligned with [31].. We end our note with some comments on it in section 6 along with possible future direction.
@@ 2 Dimensional Reduction to 2D de-Sitter in near-Nariai limit
Here we review the dimensional reduction of pure Einstein gravity in 4D to two-dimensional JT gravity in the near-Nariai limit [32],[33] following [34]. Consider the following metric ansatz of the class of spherically symmetric asymptotically 4D de Sitter metric:
@@ 2.1 Alternative dual reduction in the near-Nariai limit
Hence, the entropy is $S=S_{N}+\log(A)$. This suggests, the total entropy may have some correction which is not observed from previous way. However, this is not a useful way to study dynamics of the effective theory. This motivates us to consider an Weyl-equivalent or alternative dimensional reduction which may capture another useful and tractable description of near horizon physics in closed universe.
@@ 2.2 Partition function of massless free scalar fields in a static patch background with stretched horizon
In section (5) we will explicitly compute and will try to match this entropy using the quantization of our proposed UV free bosonic deformed CFT on a ring. We note that, the Hamiltonian of dilatons with stretched horizon boundary condition defined by some UV theory, should be identified with the total Hamiltonian of the gravitating observer plus dilatons and the observer clock is set by the ‘time’ generated by static patch Hamiltonian due to Hamiltonian constraint.
@@ 5.1 Deformed CFT on a ring, emergent Virasoro algebra and stretched horizon in static patch
Using $z=-\coth{\gamma\omega}$ and $\bar{z}=-\coth{\gamma\bar{\omega}}$, with $\omega=\theta+it^{\prime}$$\bar{\omega}=\theta-it^{\prime}$ and the transformation of stress tensor $T(\omega)=(\frac{\partial z}{\partial\omega})^{2}T(z)+\frac{c}{12}Sch\{z,\omega\}$, we end up with:
@@
Now $T(\omega)=\left[\frac{1}{\sinh^{4}{\gamma\omega}}T(z)-\frac{c\gamma^{2}}{12\pi}\right]$ and $\bar{T}(\bar{\omega})=\left[\frac{1}{\sinh^{4}{\gamma\bar{\omega}}}T(z)-\frac{c\gamma^{2}}{12\pi}\right]$. Since $\theta\rightarrow\pm\infty$ corresponds to $z\rightarrow\pm 1$, $T(z)|_{z=\pm 1}$, $\bar{T}(\bar{z})|_{\bar{z}=\pm 1}$ is finite. Thus we can immediately see [^26] Here again the stress tensor is negative but finite. Even though, one can have thermal two point function and other thermal features of the static patch, this constant negative stress tensor is not concerning. Rather the difference from an positive stress tensor in the AdS2 black hole vacuum might be quite encouraging in the sense of distinguished nature of the Nariai black hole. An observer inside the static patch can effectively measure negative local energy density.
@@ 5.2 Modular quantization on static patch and thermal partition function
[…]
  • Using this quantization, the $\epsilon\rightarrow 0$ limit can be identified with the zero modes of the Hamiltonian as we will show now. However, mapping back to $(\omega,\bar{\omega})$ plane these primaries are vanishing, leaving only vacuum at the horizon [^28] We took a different claim for these than our earlier paper [1], which we think is wrong.. Hence the vanishing edge modes should further give a vanishing micorcanonical entropy count sourced purely from the surviving vacuum state [^29] One can not obtain this just naively putting $\epsilon=0$ for which the leading part of entropy is diverging. This different results further suggests that measuring entropy without a cut-off with boundary condition is meaningless.. This on the other hand tells, the entropy explicitly measures the stretched horizon observer entropy, which can not be obtained in the semiclassical limit without an observer. The same applies in AdS2, where one must impose a cutoff on the dilatons, which can be made arbitrarily large but cannot be taken literally to infinity.
@@
Here $g_{a,b}\equiv\log{\langle 0|h\rangle\rangle}_{a,b}$ are the boundary entropies. Since in $\epsilon\rightarrow 0$ limit, $|h\rangle\rangle\rightarrow|0\rangle$, $g_{a,b}$ vanishes identically. Hence the leading term matches with the entanglement entropy in a full time slice of the static patch $\theta:(-\Lambda_{c},\Lambda_{c})$ as one can check by computing twist operator two point function using the map $z\rightarrow\omega$. This is again consistent with the fact that the time translation of static patch Hamiltonian is identified with the flow by modular Hamiltonian on the full patch bounded by the horizons.
On the other hand, we can compare this result by explicit computation of free massless bosons($c=1$) in the stretched horizon background as in (2.40). We have seen a potential mismatch of overall factor of $\frac{1}{8}$. Currently we have no clear understanding behind this mismatch. Neglecting that potential mismatch, we can see both of these terms proportional to $\log(\frac{2}{\epsilon})$ [^30] Here we took $r_{N}=1/\nu=1$. However for the general $r_{N}\neq 0$, we may need to consider $H_{c}=\alpha L_{0}+i\gamma(L_{1}-L_{-1})$ with $d=\alpha^{2}-4\gamma^{2}<0$. This class of Hamiltonian belongs to the ‘heating phase’ in the Floquet CFT language [70],[65]..
We end this section by commenting our take on the questions we have tried to address iff we may think of this deformed CFT quantization as an UV description of near Nariai black hole or extended static patch:
@@
    Thermal entropy of free massless bosons in stretched horizon background. Along the line of [65], one can again think the same as the entanglement entropy of the causal diamond vacuum with the outside [^31] This is exactly the other side of the quantization contour as described in [65].. The stretched horizons approaching the Naiai horizons or $\epsilon\rightarrow 0$ limits are the semi-classical limits in this description.
@@
    What is the (Nariai) static patch holography in this context?
    All the (dilaton) entropic degrees of freedom of the extended static patch is restored in the stretched horizon and the Hilbert space is defined by the modular quantization of worldsheet CFT.
@@
    The stretched horizon or the moving membrane defined by the worldsheet CFT with deformed Hamiltonian on a ring, plays the role of a gravitating observer, as observed by a (non-gravitating) podal observer located at the center of the static patch.
@@ 6 Discussion
In this note, our main goal is to find possible UV description of effective 2D theory of dilatons arising from reduction of higher dimensional Einstein gravity describing pure and asymptotically dS. While a stable dS2 vacuum is difficult to find in any time-dependent patches of dS2, we have managed to describe a sensible description of (extended and not pure) static patch effective theory in terms of worldsheet CFT with deformed Hamiltonian. Modular quantization of the worldsheet CFT further introduces a stretched horizon with an emergent conformal boundary condition, which acts like an internal observer in this description. In the CFT description, we have not taken $L\rightarrow iL$, since the role reversal of space-time direction in the proper time frame of the deformed Hamiltonian already manifests a CFT on 2d dS metric upto an overall change in Lorentzian signature. One may alternatively think of this set-up as de-Sitter mimicker, rather than an inherently de-Sitter. However, we would like to think of these deformed Hamiltonians of CFT generates 2d CFT on de-Sitter patches which, to the best of our knowledge, is not discussed previously in the literature. It would be definitely interesting to find a set-up where negatively divergent energy density in Milne vacuum could be compensated by gluing it with a very high positive energy density state like large AdS2 blackhole. This could probably lead to some version of cosmology in open universe.
@@
  • Source of extremal entropy from deformed CFTs: Even though, the worldsheet approach for effective dilaton theory could give us an understanding of near extremal (and near horizon) entropy in terms of entanglement entropy, the source for pure extremal entropy or Nariai entropy is not clear to us from the same worldsheet description. An external charged sector of the CFT may provide a hint towards the counting [71], [72], [73].
We end our discussion with some speculative remarks on the emergence of different worldsheet times.
@@ 6.1 Comment on emergence of observer time
Nevertheless, the main point we want to make here is that the time $t$ inherent to the Floquet Hamiltonian $H_{F}$, is a stroboscopic time $t=n(T_{1}+T_{2})$. In this Floquet sense, the stroboscopic time is emergent. To be more precise, the Floquet observer will measure stroboscopic time in their own clock which is completely different than the clocks for the observers associated to $H_{1,2}$. This is analogous to find a static patch in an accelerating or time dependent (dS) universe. Consequently, to account for the cost of residing in an accelerating universe, static observers are required to observe features like surrounding cosmological (and black hole) horizons. We know that Floquet Hamiltonians in heating phases also exhibit features such as fixed points. We feel, the gravitational observers measure the stroboscopic Floquet time in the worldsheet description, which is inherently emergent. The principle behind the fine tuning of choosing the Hamiltonians $H_{1,2}$, drive parameters and frequencies still remains elusive to us from a gravitational perspective. We hope to return with more evidences on this in near future.
Acknowledgement: The author(SD) of this paper strongly believes that any scientific progress is a collective and collaborative process where the authorship is an effort to consolidate all the ideas in a single manuscript. He would like to thank Bobby Ezhuthachan, Rahul Roy and Krishnendu Sengupta for discussion at the early stage of this project. He also wants to thank collectively Vaibhab Burman, Dibya Chakraborty, Anirban Dinda, Arnab Kundu, Parthasarathi Majumdar, Debangshu Mukherjee and Koushik Ray, as well as ChatGPT and Perplexity for related discussions and help. He is also grateful to the co-authors of [65]: Bobby Ezhuthachan, Somnath Porey and Baishali Roy for noticing and emphasizing the paper by Solodukhin [47], which provides an important clue to understand the reduction process of this note. He would also like to acknowledge the organisers of the school “Introductory School on Conformal Field Theories” at IACS(18-22 Aug,2025), which helps him to clarify some basic 2D CFT stuff. He would also like to thank the organizers as well as speakers, participants and poster presenters of Indian String Meeting (2025), who help him to clarify many doubts and conceptual mistakes regarding this work. The research work of SD is supported by a DST INSPIRE Faculty Fellowship.
@@
In [1], we claim that the classical limit of JT gravity coupled to CFT with $\mathcal{O}(1)$ central charge is described by a certain deformed CFT in a prescribed limit. Partial justification of this claim was made by the fact that in the semiclassical limit, the backreaction of the conformal fields can be neglected and the dynamics of JT coupled to matter CFT is boiled down to the dynamics of CFT in the UV frame. We recovered the entropy of the coupled system away from extremality, purely from the entropy of deformed CFT in the prescribed classical limit. Here in this appendix, we provide a concrete set-up for this justification where JT coupled to free bosons with $c=1$ arises naturally from S-wave reduction in the near extremal near horizon limit of 4D RN black holes in AdS.
@@
In either description, the Hilbert space of modular quantization or the Hilbert space of free scalars in stretched horizon background of AdS2 black hole provides the UV Hilbert space. Perhaps, the 3d version of this description along with emergent EFT description outside the horizon has been realized in [41], [103] [^44] I thank Chethan Krishnan for emphasizing this possibility at some point while discussing [41]..